OpenStudy (anonymous):

Find the measures of the numbered angles in rhombus ABCD if BE=7degrees and CD=14degrees.

Directrix (directrix):

BE=7degrees and CD=14degrees. --> These are segments and are not measured in degrees. I don't see any numbered angles. Would you check the original diagram and question and see if you wrote it correctly? Thanks.

OpenStudy (anonymous):

Let me fix it again.

Directrix (directrix):

The diagonals of a rhombus are perpendicular. A rhombus is equilateral. So, to get angle 2, use trig. cos(<2) = 7/14 = 1/2 M<2 = 60. Opposite angles of a rhombus are congruent and a diagonal of a rhombus bisects two angles of the rhombus. Putting all that together, the angles can be found. See attached diagram.

OpenStudy (anonymous):

oh you got the picture sorry my internet is working slow.